OpenStudy (anonymous):

Tina wants to save money for school. Tina invests $1000 in an account that pays an interest rate of 8%. How many years will it take for the account to reach $1300? Round your answer to the nearest hundredth.

OpenStudy (amistre64):

do we assume a compounded rate of once a year?

OpenStudy (anonymous):

I guess.

OpenStudy (anonymous):

A=P(r+1)^t

OpenStudy (amistre64):

if so, then: 1300 = 1000 (1+.08)^t ; divide out the 1000 1300/1000 = (1.08)^t ; now ln it all ln(13/10) = ln[(1.08)^t] ; apply log rules ln(13/10) = t * ln(1.08) ; and divide out the ln(1.08) ln(13/10)/ln(1.08) = t

OpenStudy (anonymous):

So it's 1.08..?

OpenStudy (amistre64):

dunno; the calculator has to figure that out ....

OpenStudy (amistre64):

google has a nice calculator feature; just type it in the search box and it should just pop out an answer

OpenStudy (anonymous):

Type in ln(13/10)/ln(1.08) = t ?

OpenStudy (amistre64):

not the "=t" part :)

OpenStudy (anonymous):

Ok thanks.

OpenStudy (amistre64):

to dbl chk the results; insert the value you get for "t" into the original equation